What mass of insulin must be dissolved in 36.0 ml of water to produce a solution with an osmotic pressure of 15.9 mmhg at 25 C? the molar mass of insulin is 5808 g/mol
Answer #1

P= 15.9 mm Hg

= (15.9/760) atm

= 0.020921 atm

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

P = C*R*T

0.020921 = C*0.0821*298.0

C =0.000855 M

volume , V = 36.0 mL

= 3.6*10^-2 L

number of mol,

n = Molarity * Volume

= 0.000855*0.036

= 3.078*10^-5 mol

mass,

m = number of mol * molar mass

= 0.000031*5808

= 0.179 g

Answer: 0.179 g